Determining A simple interest rate
you borrow $2500 froma friend and promise to pay back $2655 in six months. What simple interest rate will you pay?
example: A= P (1 + rt)
2655=2500[1 + r(0.5)]
2655=2500 + 1250r
I don't know how they got 1250r (distributive property)
Please explain
Thank you

Answer by stanbon(48551)

2655=2500[1 + r(0.5)]
2655=2500 + 1250r
I don't know how they got 1250r (distributive property)
---------
The distributive property is a*(b+c)=a*b+a*c
------------
You have: 2500*(1+r(0.5))
It becomes 2500*1 + 2500*r(0.5)
=2500 + 1250r
---------
The 1250r is (0,5)2500r = 1250r
---------
Hope that helps.
Cheers,
Stan H.